OpenStudy (anonymous):

The shop declare a discount of 20% and later a discount of 15% after first discount. Total discount was

OpenStudy (anonymous):

Price - (Price * 20/100) = Discounted price Discounted price - (discounted price * 15/100) = New Discount price Total discount = (New discount price) / (price) * 100

OpenStudy (anonymous):

Men’s while handkerchief cost Rs. 2.29 for 3. The cost per dozen is:

OpenStudy (anonymous):

set up a proportion \[ \frac{2.29 Rs}{3} = \frac{x Rs}{12}\]

OpenStudy (anonymous):

2.29/3=0.763333333 means 1 handkerchief cost Rs 0.763333333 and dozen means 12 so 0.763333333*12=9.16.

OpenStudy (anonymous):

yah. ^_^